OpenStudy (kinggeorge):

[SOLVED] Let \(\alpha := \sqrt{2-\sqrt[3]{5}} \in \mathbb{R}\) Show that \(\alpha\) is algebraic over \(\mathbb{Q}\) by finding a polynomial \(p(x)\in \mathbb{Q}[x]\) such that \(p(\alpha)=0\) Is there a better way to do this problem other than to find \(\alpha^6\) and then use guess and check with lower powers until you find the correct coefficients?

OpenStudy (zarkon):

\[p(x)=x^6-6x^4+12x^2-3\]

OpenStudy (kinggeorge):

What would be the most effective way of finding that polynomial?

OpenStudy (zarkon):

let \[x=\sqrt{2-\sqrt[3]{5}}\] \[x^2=2-\sqrt[3]{5}\] \[x^2-2=-\sqrt[3]{5}\] \[(x^2-2)^3=-5\] \[(x^2-2)^3+5=0\] \[(x^2-2)^3+5=x^6-6x^4+12x^2-3\]

OpenStudy (kinggeorge):

Awesome! Thanks.

OpenStudy (zarkon):

yw

OpenStudy (kinggeorge):

So if it were a cubed root on the outside, instead of squaring first, I would cube \(\alpha\), get all the rational numbers and \(\alpha^3\) on one side, and then cube again to the \(-\sqrt[3]{5}\) into a rational number. And when that's all done, solve for 0.

OpenStudy (zarkon):

yes

OpenStudy (kinggeorge):

Thank you.